OpenStudy (anonymous):

What is \(S_7\) for 6 – 24 + 96 – 384 + ... ?

OpenStudy (mathmale):

I'd suggest you start by determining whether or not this is a geometric series. If it is, then you'll need to identify a, the first term of the sequence, and r, the common ratio. Familiar with those?

OpenStudy (anonymous):

It is a geometric series. \[a_1 = 6\] \[6 \times -4 = -24 \times -4 = 96\] Not sure how I find the common ratio tho

OpenStudy (anonymous):

The ratio is –4. Just multiply it times each successive terms The individual terms are: 6 -24 96 -384 1536 -6144 24576 Now just add each next term to the previous partial sums S1, S1, .... and get 6 =S1 –18= –24 + 6 = S2 78 = 96 + (–18) = S3 ... etc –306 1230 –4914 19662 = S7 While this is easily done with a summation formula in one step, I think you need to go through the process manually at least once to understant just what a partial sum is.

OpenStudy (anonymous):

i think

OpenStudy (mathmale):

r is the common ratio. This means that if you multiply your first term by r, your next term will be -24. As bvb says, the common ratio is -4. It's very important that you be able to determine that your self. What I'd do would be to write 6r=-24 and then solve that for r. bvb is right in that you can answer this question by finding the 1st 7 terms of the series and then add them together for the 7th partial sum. But I urge you to do some research to find the formula for the nth partial sum of a geometric series.